PHOTOASSOCIATION OF ULTRACOLD ATOMS: A NEW TOOL FOR MOLECULAR SPECTROSCOPY

\underline{WILLIAM C. STWALLEY}, Department of Physics, University of Connecticut, Storrs, CT 06269-3046.

The field of laser cooling and trapping of atoms has had a profound impact on many areas of atomic, molecular and optical science. In particular, ultracold atoms colliding in magneto-optical traps typically have relative kinetic energies of 300μK (∼7 MHz), so free-bound absorption lines are as sharp as bound-bound lines. Moreover, because of very long range centrifugal barriers, only a few collisional partial waves are important (J′′ = 0, 1 and 2 in our work on 39K). Such photoassociative free-bound absorption allows observation of “pure long-range molecule” states (R_{e} = 28 and 39 Å in ^{39}K_{2}) as well as long-range levels of known short range states with outer classical turning points of ∼20 - 250 Å. Such spectra have been observed by “trap loss” (decrease in atomic fluorescence), molecular ionization, and fragment ionization. Optical-optical double resonance (OODR), e.g. through the “Franck-Condon windows” of the “pure long-range” states, produces readily assignable spectra at higher asymptotes as well. Further extensions of ultracold photoassociative spectroscopy and applications to determination of long range potentials, atomic properties, and Bose-Einstein condensate stability will be presented.